# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v6

@@ 7 Policy
The policy is divided into two categories: object/action-centric methods and end-to-end methods. Object/action-centric methods extract attributes from observations, such as bounding boxes, masks (Sajjan et al. (2020)), or 3D spatial action-value map (Shi et al. (2024)). These extracted attributes are then transformed into either a sequence of key poses or a single key pose, which is used in motion planning to guide robot motion. End-to-end methods directly map observation to robot action (Chi et al. (2023)). They eliminate the need for attribute extraction.
End-to-end methods are mainly divided into reinforcement learning (Herzog et al. (2023)) and imitation learning (Dasari et al. (2019)). Recent end-to-end methods have made significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
However, the above methods are all one-model-for-one-task, lacking general-purpose capability. Due to the development of foundation models, general-purpose models have advanced. The representation of task instruction can be categorized into four types: language, human video, goal image, and multimodal prompts. BC-Z (Jang et al. (2022)) and Vid2Robot (Jain et al. (2024)) introduce a video-conditioned policy that uses human video as task instructions. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates actions for manipulation by combining the target image with the observation image. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023c)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. MUTEX (Shah et al. (2023)) extends instruction to various modalities and develops speech-conditioned, speech-goal-conditioned, image-goal-conditioned, and text-goal-conditioned.
Language-conditioned general-purpose policies remain the predominant paradigm in current research. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Key-Pose (VLAKP), Vision-Language-Action-Dense-Pose (VLADP). VLAKP is more similar to traditional object/action-centric approaches. On the other hand, VLADP resembles traditional end-to-end methods. Recent studies have explored the use of foundation models to directly synthesize low-level policy code (Liang et al. (2023); Yoshida et al. (2025)). Such policies generate executable code for robots, enabling fine-grained human inspection and debugging. We denote this paradigm as Vision-Language-Action-Code (VLAC).
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (Trivedi et al. (2021). However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (Ellis et al. (2023)). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (Chen et al. (2021b); Austin et al. (2021)). Code-As-Policy (Liang et al. (2023)) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. From-text-to-motion (Yoshida et al. (2025)) translates descriptions of human actions into humanoid robot motion code, enabling it to perform various tasks autonomously.
@@ 7.2 VLAKP
Utilizing foundation models pre-trained on existing large-scale internet datasets enables the direct perception of observation images and outputting key poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within an observation image and outputs the 3D position of these manipulated objects. Voxposer (Huang et al. (2023c)) utilizes LLMs to generate code that interacts with VLMs to produce affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation. ReKep (Huang et al. (2024a)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) demonstrates the capability of imitation learning in language-conditioned general manipulation. However, CLIPort (Shridhar et al. (2021)) addresses 4-DoF end-effector pose prediction by treating it as a pixel classification problem. Keypoint-based approaches are extended to handle 6-DoF end-effector pose prediction. Due to keypoint-based methods primarily focus on 3D scene-to-action tasks, these methods become computationally expensive as resolution requirements increase (Ke et al. (2024)). To address high spatial resolution, PerAct (Shridhar et al. (2023)) uses the latent set self-attention of Perceiver (Jaegle et al. (2021)), which has linear complexity with voxels. Act3D (Gervet et al. (2023)) represents scenes as a continuous 3D feature field, transforming 2D model features into 3D feature clouds using sensed depth and learns a 3D feature field of arbitrary spatial resolution through recurrent coarse-to-fine point sampling.
@@
Some research has extended the work on PerAct (Shridhar et al. (2023)) and Act3D (Gervet et al. (2023)). ChainedDiffuser (Xian et al. (2023)) builds upon Act3D (Gervet et al. (2023)) by replacing the motion planner with a diffusion model. This approach addresses the challenges of continuous interaction tasks. The 3D Diffuser Actor (Ke et al. (2024)), similar to Act3D (Gervet et al. (2023)), employs tokenized 3D scene representations. However, unlike Act3D (Gervet et al. (2023)) and 3D Diffusion Policy (Ze et al. (2024b)) with 1D point cloud embeddings, 3D Diffuser Actor (Ke et al. (2024)) leverages CLIP to extract features from 2D images and aggregates them into a 3D scene representation. GNFactor (Ze et al. (2023)) improves upon PerAct (Shridhar et al. (2023)) by enhancing 3D semantic features. It achieves this by distilling pre-trained semantic features from 2D foundation models into Neural Radiance Fields (NeRFs). DNAct (Yan et al. (2024)) builds on GNFactor (Ze et al. (2023)) and transforms the perceiver model into a diffusion head. VoxAct-B (Liu et al. (2024b)) uses VLM to divide the task into subtasks for the left arm and the right arm and applies PerAct (Shridhar et al. (2023)) to generate separate key poses for each arm.
@@
The current imitation learning approaches also include methods using large-scale LLM/VLM. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. Xu et al. (2024) considers the object motion produced by LLM/VLM, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the key pose. However, the key pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point. 3D-VLA (Zhen et al. (2024)) can generate the final state image and point cloud based on user input. This goal state is then used to create key poses in 3D VLA.
@@
Figure 7: Foundation Models for Policy. VLAC integrates task instruction, pre-written APIs, and example inputs into LLM. It generates corresponding execution code (Liang et al. (2023); Yoshida et al. (2025)). 2) VLAKP without foundation models training inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation or keypoints, and both the manipulated object and object feature images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D key pose (Huang et al. (2023b)) or relational keypoint constraints (Huang et al. (2024a)). 3) VLAKP with training uses LLM to extract tokens from Task Instructions. VFM or VLM extract features from Multi-View Images and generate 3D features through 3D Lift. Finally, these features, along with proprioception, are used as input to the Diffusion Model to generate the key pose (Ke et al. (2024)). 4) VLADP with training (Single-System) outputs dense pose and the hindsight image by inputting task instruction and observation into a pre-trained model after training (Black et al. ). The difference between VLADP and VLAKP with training lies in generating dense poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to key pose. Key pose often requires subsequent motion planning. Outputting dense pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining after the next movement, predicting both the next action and the hindsight image can improve the performance (Bousmalis et al. (2023)). 5) The VLADP with training (Double-System) generates dense pose by using models with different inference frequencies (Figure.ai (2025)). This method can effectively leverage prior knowledge from VLM and improve inference frequency. 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (Di Palo et al. (2023)), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (Ma et al. (2023b)), while VLM can utilize prior knowledge to provide predicted action and sparse/dense reward, enhancing the effective exploration in reinforcement learning.(Ye et al. (2023a)).
@@
VLAC There are many Internet code datasets available. These can serve as priors for code generation. This helps improve the generalization of VLAC-style methods (Liang et al. (2023)). VLAC facilitates human attention to fine-grained details, enhancing the debugging process. VLAC is able to generate more creative actions, as shown in Yoshida et al. (2025). The use of third-party libraries in VLAC can constrain the extensibility of policy tasks. Syntax errors and similar simple faults tend to appear frequently..
VLAKP Compared to VLADP, VLAKP has a key advantage. It can leverage rich prior knowledge from foundation models in CV and NLP. This supports object-centric and action-centric action generation. As a result, VLAKP requires less training data and generalizes better than VLADP (Ke et al. (2024)). VLAKP improves upon VLAC by replacing sequential chaining with spatial composition under joint optimization. This flexibility supports more diverse manipulation tasks and ensures more stable execution (Huang et al. (2023c)). Mostly relies on 3D inputs, sensitive to pose errors. Requires precise scene reconstruction and calibration. VLAKP often requires motion planning. However, motion planning is not always reliable.
@@ 7.3 VLADP
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (Reed et al. (2022)) and RT-1 (Brohan et al. (2022)) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Building upon Gato, RoboCat (Bousmalis et al. (2023)) demonstrates that a large sequence model can learn unseen tasks through few-shot learning. It proposes a simple but effective self-improvement process. Additionally, it shows that predicting both the next action and the hindsight image after executing that action can enhance performance. Building upon RT-1, RoboAgent (Bharadhwaj et al. (2023)) enhances model generalization and stability through data augmentation and action-chunking. MOO (Stone et al. (2023)) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
@@
As for multi-task reinforcement learning. PI-QT-Opt (Lee et al. (2023)) leverages a large-scale, multi-task dataset and employs a model-free off-policy reinforcement learning approach for training. Q-Transformer (Chebotar et al. (2023)) facilitates training high-capacity sequential architectures on mixed-quality data by applying transformer models to RL.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024b)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating pose as tokens. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023d)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
@@
However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)) and HPT (Wang et al. (2024d)) propose multi-module networks to address this issue. RDT-1B (Liu et al. (2024d)) and PI (Black et al. (2024)) propose unified action space to address this issue. The Open X-Embodiment dataset lacks in-the-wild scenes. DROID (Khazatsky et al. (2024)) introduces an “in-the-wild” robot manipulation dataset. It contains 76k trajectories, or 350 hours of interaction data. This data is collected from 564 scenes, 86 tasks, and 52 buildings over 12 months.
@@
Internet videos contain information on the physics and dynamics of the world, some studies have explored training foundation models using both video datasets and manipulation data. SuSIE (Black et al. (2023)) uses an image-editing diffusion model. This model is fine-tuned on human videos and robot rollouts. It acts as a high-level planner by proposing intermediate subgoals. These subgoals can be accomplished by a low-level controller. The two training steps of SuSIE do not share weights. GR-1 (Wu et al. (2023a)) is initially trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. GR-2 (Cheang et al. (2024)) uses VQGAN to convert each image into discrete tokens and is trained with a larger text-video dataset than GR-1. LAPA (Ye et al. (2024)) begins by extracting the latent delta action between video frames. It then labels the video dataset with this information. These labeled datasets are used to train a VLM network. Finally, a small-scale robot manipulation dataset is applied for fine-tuning, enabling the mapping of latent actions to robot actions. Go-1 (Bu et al. (2025)) trains latent actions using human videos and combines those with manipulation data. Then, it trains the VLA on the merged dataset to boost the model’s generalization.
@@
Previous studies, such as GR-1 (Wu et al. (2023a)) and GR-2 (Cheang et al. (2024)), train the policy head using MSE regression. In contrast, OpenVLA (Kim et al. (2024)) and RT-2 (Brohan et al. (2023)) apply next-token prediction for their policy head. Building on the success of diffusion policy (Chi et al. (2023)), PI0 (Black et al. ) and TinyVLA (Wen et al. (2024b)) adopt diffusion head as its policy head, achieving better performance than OpenVLA. In order to address higher degree of multi-modality in the distribution of feasible actions for bimanual manipulation, RDT-1B (Liu et al. (2024d)) utilizes Diffusion Transformers (DiTs) as its scalable backbone network.
@@
Although the diffusion policy can represent complex continuous action distributions, OpenVLA-OFT (Kim et al. (2025)) has shown in dual-arm tasks that fine-tuning the VLA with an L1 regression objective achieves performance similar to diffusion-based fine-tuning. However, it offers faster training convergence and inference speed. FAST (Pertsch et al. (2025)) proposes a new compression-based tokenization scheme for next-token prediction. This method matches the performance of diffusion VLAs, while reducing training time by up to 5x across multiple dexterous manipulation tasks.
@@
Previous methods face a basic tradeoff: VLM backbones are general but slow, while robot visuomotor policies are fast but not general. Synchronizing both does not improve inference speed. Helix (Figure.ai (2025)) and Groot N1 (Bjorck et al. (2025)) overcome this tradeoff with two asynchronous complementary systems, trained end-to-end to communicate. However, Groot N1 makes more use of human video latent action and simulated data compared to Helix.

### v7

@@ 7 Policy
The policy is divided into two categories: object/action-centric methods and end-to-end methods. Object/action-centric methods extract attributes from observations, such as bounding boxes, masks (sajjan2020clear), or 3D spatial action-value map (shi2024asgrasp). These extracted attributes are then transformed into either a sequence of key poses or a single key pose, which is used in motion planning to guide robot motion. End-to-end methods directly map observation to robot action (chi2023diffusion). They eliminate the need for attribute extraction.
End-to-end methods are mainly divided into reinforcement learning (herzog2023deep) and imitation learning (dasari2019robonet). Recent end-to-end methods have made significant progress. ACT (zhao2023learning) uses action chunks to reduce compounding errors. Diffusion policy (chi2023diffusion) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
However, the above methods are all one-model-for-one-task, lacking general-purpose capability. Due to the development of foundation models, general-purpose models have advanced. The representation of task instruction can be categorized into four types: language, human video, goal image, and multimodal prompts. BC-Z (jang2022bc) and Vid2Robot (jain2024vid2robot) introduce a video-conditioned policy that uses human video as task instructions. DALL-E-Bot (Kapelyukh_2023) employs DALL-E to generate target images for tasks and generates actions for manipulation by combining the target image with the observation image. VIMA (jiang2023vima) and MIDAS (li2023mastering) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. MUTEX (shah2023mutex) extends instruction to various modalities and develops speech-conditioned, speech-goal-conditioned, image-goal-conditioned, and text-goal-conditioned.
Language-conditioned general-purpose policies remain the predominant paradigm in current research. RT-2 (rt2) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Key-Pose (VLAKP), Vision-Language-Action-Dense-Pose (VLADP). VLAKP is more similar to traditional object/action-centric approaches. On the other hand, VLADP resembles traditional end-to-end methods. Recent studies have explored the use of foundation models to directly synthesize low-level policy code (liang2023code; yoshida2025text). Such policies generate executable code for robots, enabling fine-grained human inspection and debugging. We denote this paradigm as Vision-Language-Action-Code (VLAC).
@@ 7.1 VLAC
Code generation and program synthesis have been demonstrated to be capable of developing generalizable, interpretable policy (trivedi2021learning. However, a robot capable of generating code for multiple tasks, rich knowledge across various domains is essential (ellis2023dreamcoder). Therefore, scholars aim to apply the prior knowledge of LLM to code generation task (chen2021evaluating; austin2021program). Code-As-Policy (liang2023code) demonstrates the possibility of using LLMs to directly generate code for robot execution based on prompts. The study shows that 1) code-writing LLMs enable novel reasoning capability, such as encoding spatial relationships by leveraging familiarity with third-party libraries and 2) hierarchical code-writing inspired by recursive summarization improves code generation. From-text-to-motion (yoshida2025text) translates descriptions of human actions into humanoid robot motion code, enabling it to perform various tasks autonomously.
@@ 7.2 VLAKP
Utilizing foundation models pre-trained on existing large-scale internet datasets enables the direct perception of observation images and outputting key poses. Instruct2Act (huang2023instruct2act) utilizes CLIP and SAM to identify manipulated objects within an observation image and outputs the 3D position of these manipulated objects. Voxposer (huang2023voxposer) utilizes LLMs to generate code that interacts with VLMs to produce affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation. ReKep (huang2024rekep) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
As for imitation learning methods, CLIPort (shridhar2021cliport) demonstrates the capability of imitation learning in language-conditioned general manipulation. However, CLIPort (shridhar2021cliport) addresses 4-DoF end-effector pose prediction by treating it as a pixel classification problem. Keypoint-based approaches are extended to handle 6-DoF end-effector pose prediction. Due to keypoint-based methods primarily focus on 3D scene-to-action tasks, these methods become computationally expensive as resolution requirements increase (ke20243d). To address high spatial resolution, PerAct (shridhar2023perceiver) uses the latent set self-attention of Perceiver (jaegle2021perceiver), which has linear complexity with voxels. Act3D (gervet2023act3d) represents scenes as a continuous 3D feature field, transforming 2D model features into 3D feature clouds using sensed depth and learns a 3D feature field of arbitrary spatial resolution through recurrent coarse-to-fine point sampling.
@@
Some research has extended the work on PerAct (shridhar2023perceiver) and Act3D (gervet2023act3d). ChainedDiffuser (xian2023chaineddiffuser) builds upon Act3D (gervet2023act3d) by replacing the motion planner with a diffusion model. This approach addresses the challenges of continuous interaction tasks. The 3D Diffuser Actor (ke20243d), similar to Act3D (gervet2023act3d), employs tokenized 3D scene representations. However, unlike Act3D (gervet2023act3d) and 3D Diffusion Policy (ze20243d) with 1D point cloud embeddings, 3D Diffuser Actor (ke20243d) leverages CLIP to extract features from 2D images and aggregates them into a 3D scene representation. GNFactor (ze2023gnfactor) improves upon PerAct (shridhar2023perceiver) by enhancing 3D semantic features. It achieves this by distilling pre-trained semantic features from 2D foundation models into Neural Radiance Fields (NeRFs). DNAct (yan2024dnact) builds on GNFactor (ze2023gnfactor) and transforms the perceiver model into a diffusion head. VoxAct-B (liu2024voxact) uses VLM to divide the task into subtasks for the left arm and the right arm and applies PerAct (shridhar2023perceiver) to generate separate key poses for each arm.
@@
The current imitation learning approaches also include methods using large-scale LLM/VLM. LEO (huang2023embodied) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. xu2024manifoundation considers the object motion produced by LLM/VLM, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the key pose. However, the key pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point. 3D-VLA (zhen20243d) can generate the final state image and point cloud based on user input. This goal state is then used to create key poses in 3D VLA.
@@
Figure 7: Foundation Models for Policy. VLAC integrates task instruction, pre-written APIs, and example inputs into LLM. It generates corresponding execution code (liang2023code; yoshida2025text). 2) VLAKP without foundation models training inputs task instruction into LLM, which specifies the manipulated object. The observation image is fed into VFM for object segmentation or keypoints, and both the manipulated object and object feature images are input into VLM, outputting the pixel mapping of the object to be manipulated into cartesian 3D key pose (huang2023instruct2act) or relational keypoint constraints (huang2024rekep). 3) VLAKP with training uses LLM to extract tokens from Task Instructions. VFM or VLM extract features from Multi-View Images and generate 3D features through 3D Lift. Finally, these features, along with proprioception, are used as input to the Diffusion Model to generate the key pose (ke20243d). 4) VLADP with training (Single-System) outputs dense pose and the hindsight image by inputting task instruction and observation into a pre-trained model after training (black2410pi0). The difference between VLADP and VLAKP with training lies in generating dense poses directly through policies, allowing for conversion into trajectory through time sequences, offering a more end-to-end approach compared to key pose. Key pose often requires subsequent motion planning. Outputting dense pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. However, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layer. As for imagining after the next movement, predicting both the next action and the hindsight image can improve the performance (bousmalis2023robocat). 5) The VLADP with training (Double-System) generates dense pose by using models with different inference frequencies (helix2025). This method can effectively leverage prior knowledge from VLM and improve inference frequency. 6)Foundation Models assisting for Reinforment Learning. LLM generates subgoals based on task instruction to transform long-horizon tasks into short-horizon ones (di2023towards), facilitating RL learning. LLM also creates a reward function for RL according to task instruction (ma2023eureka), while VLM can utilize prior knowledge to provide predicted action and sparse/dense reward, enhancing the effective exploration in reinforcement learning.(ye2023foundation).
@@
VLAC There are many Internet code datasets available. These can serve as priors for code generation. This helps improve the generalization of VLAC-style methods (liang2023code). VLAC facilitates human attention to fine-grained details, enhancing the debugging process. VLAC is able to generate more creative actions, as shown in yoshida2025text. The use of third-party libraries in VLAC can constrain the extensibility of policy tasks. Syntax errors and similar simple faults tend to appear frequently..
VLAKP Compared to VLADP, VLAKP has a key advantage. It can leverage rich prior knowledge from foundation models in CV and NLP. This supports object-centric and action-centric action generation. As a result, VLAKP requires less training data and generalizes better than VLADP (ke20243d). VLAKP improves upon VLAC by replacing sequential chaining with spatial composition under joint optimization. This flexibility supports more diverse manipulation tasks and ensures more stable execution (huang2023voxposer). Mostly relies on 3D inputs, sensitive to pose errors. Requires precise scene reconstruction and calibration. VLAKP often requires motion planning. However, motion planning is not always reliable.
@@ 7.3 VLADP
Effective robotic multi-task learning necessitates a high-capacity model, hence Gato (reed2022generalist) and RT-1 (rt1) devise transformer-based architectures. Nonetheless, RT-1 and Gato differ; RT-1’s input lacks proprioception from the robot body, while Gato incorporates proprioception. Building upon Gato, RoboCat (bousmalis2023robocat) demonstrates that a large sequence model can learn unseen tasks through few-shot learning. It proposes a simple but effective self-improvement process. Additionally, it shows that predicting both the next action and the hindsight image after executing that action can enhance performance. Building upon RT-1, RoboAgent (bharadhwaj2023roboagent) enhances model generalization and stability through data augmentation and action-chunking. MOO (stone2023open) leverages Owl-ViT to extract object locations from observation images, enhancing RT-1’s open-set detection capability.
@@
As for multi-task reinforcement learning. PI-QT-Opt (lee2023pi) leverages a large-scale, multi-task dataset and employs a model-free off-policy reinforcement learning approach for training. Q-Transformer (q-transformer) facilitates training high-capacity sequential architectures on mixed-quality data by applying transformer models to RL.
@@
Utilizing pre-trained VLMs (zhang2024vision) for fine-tuning to construct RFMs is considered efficient. RT-2 (rt2) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (chen2023pali) and PaLM-E (palme) after treating pose as tokens. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (li2023vision) and OpenVLA (kim2024openvla) emerge, effectively enabling a robot manipulation policy with VLMs.
@@
However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (RTX) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (team2024octo) and HPT (wang2024scaling) propose multi-module networks to address this issue. RDT-1B (liu2024rdt) and PI (black2024pi_0) propose unified action space to address this issue. The Open X-Embodiment dataset lacks in-the-wild scenes. DROID (khazatsky2024droid) introduces an “in-the-wild” robot manipulation dataset. It contains 76k trajectories, or 350 hours of interaction data. This data is collected from 564 scenes, 86 tasks, and 52 buildings over 12 months.
@@
Internet videos contain information on the physics and dynamics of the world, some studies have explored training foundation models using both video datasets and manipulation data. SuSIE (black2023zero) uses an image-editing diffusion model. This model is fine-tuned on human videos and robot rollouts. It acts as a high-level planner by proposing intermediate subgoals. These subgoals can be accomplished by a low-level controller. The two training steps of SuSIE do not share weights. GR-1 (wu2023unleashing) is initially trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. GR-2 (cheang2024gr) uses VQGAN to convert each image into discrete tokens and is trained with a larger text-video dataset than GR-1. LAPA (ye2024latent) begins by extracting the latent delta action between video frames. It then labels the video dataset with this information. These labeled datasets are used to train a VLM network. Finally, a small-scale robot manipulation dataset is applied for fine-tuning, enabling the mapping of latent actions to robot actions. Go-1 (bu2025agibot) trains latent actions using human videos and combines those with manipulation data. Then, it trains the VLA on the merged dataset to boost the model’s generalization.
@@
Previous studies, such as GR-1 (wu2023unleashing) and GR-2 (cheang2024gr), train the policy head using MSE regression. In contrast, OpenVLA (kim2024openvla) and RT-2 (rt2) apply next-token prediction for their policy head. Building on the success of diffusion policy (chi2023diffusion), PI0 (black2410pi0) and TinyVLA (wen2024tinyvla) adopt diffusion head as its policy head, achieving better performance than OpenVLA. In order to address higher degree of multi-modality in the distribution of feasible actions for bimanual manipulation, RDT-1B (liu2024rdt) utilizes Diffusion Transformers (DiTs) as its scalable backbone network.
@@
Although the diffusion policy can represent complex continuous action distributions, OpenVLA-OFT (kim2025fine) has shown in dual-arm tasks that fine-tuning the VLA with an L1 regression objective achieves performance similar to diffusion-based fine-tuning. However, it offers faster training convergence and inference speed. FAST (pertsch2025fast) proposes a new compression-based tokenization scheme for next-token prediction. This method matches the performance of diffusion VLAs, while reducing training time by up to 5x across multiple dexterous manipulation tasks.
@@
Previous methods face a basic tradeoff: VLM backbones are general but slow, while robot visuomotor policies are fast but not general. Synchronizing both does not improve inference speed. Helix (helix2025) and Groot N1 (bjorck2025gr00t) overcome this tradeoff with two asynchronous complementary systems, trained end-to-end to communicate. However, Groot N1 makes more use of human video latent action and simulated data compared to Helix.